# Topological transition in filamentous cyanobacteria: from motion to structure

### v2

@@ IV Conclusions
We have studied the dynamics and structural properties of a colony of filamentous cyanobacteria as they undergo a topological transition to a reticulate pattern. Their ability to glide over one another while experiencing weak aligning interactions enables the formation of a disordered network-like structure of filaments within bundles that link together. In addition to the network topology, confined cyanobacteria systems can form topological defects [76, 75], bearing a striking resemblance to some active nematic systems [97, 98]. Other than these confined systems, where defects can only resolve the mismatch of order by non-local reconfiguration, the unconfined cyanobacteria can glide over one another, resulting in dynamics not dominated by topological defects.
@@
Filamentous cyanobacteria, which we have termed active spaghetti [72], constitute a new class of active matter, due to the path-tracking nature of their gliding motility. We derived the nondimensional equations for a minimal model of their behavior that is firmly grounded in experimental evidence. The model captures the transient dynamics leading to the formation of the reticulate pattern. It also predicts a crossover in the mean-square displacement of filaments between ballistic and diffusive regimes, which can be rationalized in terms of an effective diffusivity. We explored the dynamics for different filament lengths, and found that the order parameter appears to approach a second order or continuous phase transition in the limit of very short filaments. This is in contrast to the well-studied Vicsek model [16] which may exhibit a weak-first order transition [99]. Furthermore, our system does not exhibit an ordered phase where all filaments point in the same direction, on account of the curvature fluctuations present in our model. The Péclet number dictates the typical length scale of the patterns, and together with the area coverage $\Phi$ describes the relevant fields of the nonequilibrium phase diagram for our system. Inspired by polymer physics, we introduced parallel and perpendicular pair correlation functions, which capture the fundamental structural properties of the reticulate bundles. In particular, the minimum of $g_{\perp}(r)$ faithfully predicts the typical length scale $\ell^{*}$ of the pattern. These observations can help further understand the formation and properties of cyanobacterial biomats.
@@
Figure 11: Mature cyanobacteria biofilm. O. lutea grown on agar for six months shows a more complex, three-dimensional structure. Photographs of the top surface of the biomat (left panel: face-on view; right panel: low-angle shot of the same sample) shows a complex topography of raised, filled circular features, which develop out of an earlier template of a more open, reticulate structure.
@@
Biofilms and biomats are self-organized microbial communities that differentiate and develop phenotypic adaptation both in space and time [8]. It stands to reason that the form of the reticulate patterns studied here has a biological function sculpted by survival needs. The reticulate network then serves as a foundation for mature biofilms that adapts to the features of the surrounding environment. Even growing as films on glass culture bottles, these biofilms can easily reach 100 $\mu$m thick, much thicker than an individual filament. Because gliding cyanobacteria are surface associated, a natural question is then to inquire how the surface properties influence the growth and maturation of the biofilm [100, 101]. As an initial exploration of the effects of substrate stiffness on colony formation, we also cultivated O. lutea on agar plates, see Fig. 11. These were prepared by mixing 1.5% agar, by weight, with a 1:50 dilution of BG11 broth in deionized water, following the protocols of Ref. [102]. Every eight weeks, 2 ml of 1:100 BG11 broth were added to the plates. Samples were otherwise left undisturbed, growing under identical conditions to those described for the transient experiment (Fig. 2), and following Ref. [72]. Over time, the initially reticulate pattern developed into a three-dimensional structure, with the open areas of the network filling in and growing upward that was templated on the original network boundaries. These observations point at the physical richness of this biological system and invite future exploration.

### v3

@@ IV Conclusions
We have studied the dynamics and structural properties of colonies of filamentous cyanobacteria as they undergo a topological transition to a reticulate pattern. The ability of cyanobacteria to glide over one another while experiencing weak aligning interactions enables the formation of a disordered network-like structure of filaments within bundles that link together. In addition to the network topology, confined cyanobacteria systems can form topological defects [76, 75], bearing a striking resemblance to some active nematic systems [98, 99]. Other than these confined systems, where defects can only resolve the mismatch of order by non-local reconfiguration, the unconfined cyanobacteria can glide over one another, resulting in dynamics not dominated by topological defects.
@@
Filamentous cyanobacteria, which we have termed active spaghetti [72], constitute a new class of active matter, due to the path-tracking nature of their gliding motility. We derived the nondimensional equations for a minimal model of their behavior that is firmly grounded in experimental evidence. The model captures the transient dynamics leading to the formation of the reticulate pattern. It also predicts a crossover in the mean-square displacement of filaments between ballistic and diffusive regimes, which can be rationalized in terms of an effective diffusivity. We explored the dynamics for different filament lengths, and found that the order parameter appears to approach a continuous phase transition in the limit of very short filaments. This is in contrast to the well-studied Vicsek model [16] which exhibits a weak first-order transition [100]. Furthermore, our system does not exhibit an ordered phase where all filaments point in the same direction, on account of the curvature fluctuations present in our model. The Péclet number dictates the typical length scale of the patterns, and together with the area coverage $\Phi$ describes the relevant fields of the nonequilibrium phase diagram for our system. Inspired by polymer physics, we introduced parallel and perpendicular pair correlation functions, which capture the fundamental structural properties of the reticulate bundles. In particular, the minimum of $g_{\perp}(r)$ faithfully predicts the typical length scale $\ell^{*}$ of the pattern. These observations can help further understand the formation and properties of cyanobacterial biomats.
@@
Figure 11: Mature cyanobacteria biofilm. O. lutea grown on agar for six months shows a more complex, three-dimensional structure. Photographs of the top surface of the biomat (left panel: face-on view; right panel: low-angle shot of the same sample) show a complex topography of raised, filled circular features, which develop out of an earlier template of a more open, reticulate structure.
@@
Biofilms and biomats are self-organized microbial communities that differentiate and develop phenotypic adaptation both in space and time [8]. It stands to reason that the form of the reticulate patterns studied here has a biological function sculpted by survival needs. The reticulate network then serves as a foundation for mature biofilms that adapts to the features of the surrounding environment. Even growing as films on glass culture bottles, these biofilms can easily reach 100 $\mu$m thick, much thicker than an individual filament. Because gliding cyanobacteria are surface associated, a natural question is then to inquire how the surface properties influence the growth and maturation of the biofilm [101, 102]. As an initial exploration of the effects of substrate stiffness on colony formation, we also cultivated O. lutea on agar plates, see Fig. 11. These were prepared by mixing 1.5% agar, by weight, with a 1:50 dilution of BG11 broth in deionized water, following the protocols of Ref. [103]. Every eight weeks, 2 ml of 1:100 BG11 broth were added to the plates. Samples were otherwise left undisturbed, growing under identical conditions to those described for the transient experiment (Fig. 2), and following Ref. [72]. Over time, the initially reticulate pattern developed into a three-dimensional structure, with the open areas of the network filling in and growing upward, which was templated on the original network boundaries. These observations point at the physical richness of this biological system and invite future exploration.